OpenStudy (yanasidlinskiy):

How in the world am I supposed to solve for x? a(bx+2)=cx-12

jagr2713 (jagr2713):

hold on

OpenStudy (oleg3321):

@jagr2713 please help me with my problem

jagr2713 (jagr2713):

Divide each term in the equation by (bx+2). Divide each term in the equation by (bx+2). a(bx+2)/bx+2=cx/bx+2−12/bx+2

jagr2713 (jagr2713):

Reduce the expression by canceling out the common factor of (bx+2) from the numerator and denominator. a (bx+2)/ (bx+2) =cx/bx+2−12/bx+2 Reduce the expression by canceling out the common factor of (bx+2) from the numerator and denominator. a=cx/bx+2−12/bx+2 Combine all similar expressions in the polynomial. a=cx−12bx+2

jagr2713 (jagr2713):

you get it?

OpenStudy (yanasidlinskiy):

You know I'm solving for x. Right?

OpenStudy (yanasidlinskiy):

Cuz, it doesn't seem like you solved for x.

jagr2713 (jagr2713):

lol i sloved for a hold on let me type it

jagr2713 (jagr2713):

its long can i just give answer?

jagr2713 (jagr2713):

Multiply a by each term inside the parentheses. Multiply a by each term inside the parentheses (bx+2). a(bx)+a(2)=cx−12 Multiply a by the bx inside the parentheses. a⋅bx+a(2)=cx−12 Multiply a by bx to get abx. abx+a(2)=cx−12 Multiply a by the 2 inside the parentheses. abx+a⋅2=cx−12 Multiply a by 2 to get 2a. abx+2a=cx−12

jagr2713 (jagr2713):

Since cx contains the variable to solve for, move it to the left-hand side of the equation by subtracting cx from both sides. abx+2a−cx=−12 Factor out the GCF of x from each term in the polynomial.

jagr2713 (jagr2713):

Factor out the GCF of x from the expression abx. x(ab)−cx=−2a−12 Factor out the GCF of x from the expression −cx. x(ab)+x(−c)=−2a−12 x(ab)+x(−c)=−2a−12 Factor out the GCF of x from abx−cx. x(ab−c)=−2a−12 Divide each term in the equation by (ab−c).

jagr2713 (jagr2713):

divide each term in the equation by (ab−c). x(ab−c)/ab−c=−2a/ab−c−12/ab−c Reduce the expression by canceling out the common factor of (ab−c) from the numerator and denominator. x (ab−c) /(ab−c) =−2a/ab−c−12/ab−c Reduce the expression by canceling out the common factor of (ab−c) from the numerator and denominator. x=−2a/ab−c−12/ab−c Combine all similar expressions in the polynomial. x=−2/a−12/ab−c Factor out the GCF of −2 from the expression −2a. x=−2(a)−12/ab−c Factor out the GCF of −2 from the expression −12. x=−2(a)−2(6/)ab−c Factor out the GCF of −2 from −2a−12. x=−2(a+6)/ab−c Remove the parentheses from the numerator. x=−2(a+6/)ab−c

jagr2713 (jagr2713):

you get it

OpenStudy (yanasidlinskiy):

I don't have any of that as my answer choice. Here's what it is: A. x=(-12-2a)/(ab-c);ab≠c B. x=(-12+2a)/(c+ab);ab≠c C. x=(2a+12)/(c-ab);no restrictions D. x=(ab-c)/(-12-2a);a≠-1

jagr2713 (jagr2713):

wow idk sorry

OpenStudy (zzr0ck3r):

a(bx+2)=cx-12 abx+2a=cx-12 abx-cx=-2a-12 x(ab-c)=-2a-12 x=(-2a-12)/(ab-c)

OpenStudy (yanasidlinskiy):

It's ok, jagr. At least you tried:D Gave it a shot:D

OpenStudy (zzr0ck3r):

=(-12-2a)/(ab-c)

OpenStudy (anteater):

I think it should be A The answer you got was: x=−2(a+6/)ab−c That would be (-2a-12)/(ab-c) if you distribute the 2; then change the order of the terms in the numerator: (-12-2a)/(ab -c) because that would be an equivalent expression

OpenStudy (anteater):

And, of course ab can't equal c, because if it did you would be dividing by 0 ;)

OpenStudy (yanasidlinskiy):

Ok. So what is it? I need an explanation. I don't really understand this. Everyone has different answers and I don't even have a clue who to go with.

jagr2713 (jagr2713):

this is an confusing question lol

OpenStudy (anteater):

Jagr's answer is correct. However, it doesn't "appear" to show up in your list of choices - but actually it does.

OpenStudy (anteater):

Here is why. You got the answer x = -2(a + 6)/(ab - c)

OpenStudy (anteater):

If you distribute the -2 in the numerator, you get the equivalent expression (-2a -12)/(ab-c)

OpenStudy (anteater):

You can then use the commutative property of addition on the numerator as well, to get another equivalent expression: (-12-2a)/(ab-c)

OpenStudy (anteater):

So that answer is actually the "same" as the one you got, it just looks a bit different!

OpenStudy (yanasidlinskiy):

Lol, Let me ask @ganeshie8 . Lets see what he says.

OpenStudy (anteater):

Sure! :D

OpenStudy (zzr0ck3r):

@YanaSidlinskiy what do you not understand about my solution its only a few steps so let me know which part confuses you

OpenStudy (zzr0ck3r):

\(a(bx+2)=cx-12 \\abx+2a=cx-12 \\abx-cx=-2a-12 \\x(ab-c)=-2a-12 \\x=\frac{-12-2a}{ab-c}\)

OpenStudy (yanasidlinskiy):

Why are you subtracting cx in the third step? How come you moved all the numbers to the right and variables to the left?

OpenStudy (zzr0ck3r):

we want to get all the terms with x on one sides and all the other terms on the other, so that we can isolate x

OpenStudy (yanasidlinskiy):

Oh......Ok..Thank you soooooooooo much:D

OpenStudy (zzr0ck3r):

np

jagr2713 (jagr2713):

Nice zz

OpenStudy (yanasidlinskiy):

Oh, shoot! Forgot to say \(\huge\cal\color{lime}{Thank~you!!!}\) to everyone who replied and tried to help me out!!:D Thanx again!!:D

jagr2713 (jagr2713):

yw lol

OpenStudy (anteater):

Wlcm :) Have a good afternoon!

OpenStudy (yanasidlinskiy):

Hahha:D Thanx anteater:D you too:D